OpenStudy (anonymous):

Solve for x. 3(3x - 1) + 2(3 - x) = 0 x = 3/7 x = -3/7 x = 1/3 x = -1/3

OpenStudy (anonymous):

3(3x - 1) + 2(3 - x) = 0 9x - 3 + 6 - 2x =0 whats next

OpenStudy (anonymous):

combine like terms!! :)

OpenStudy (anonymous):

8x = -3 | : 8 x = -(3/8) correct me if I'm wrong :)

OpenStudy (anonymous):

@addinol you are wrong

OpenStudy (anonymous):

help her walk through the problem

OpenStudy (anonymous):

9x - 3 + 6 - 2x =0 7x + 3 =0 7x = -3 x = -3/7

OpenStudy (farmdawgnation):

@addinol Two things. 1. Please don't interrupt other users trying to work with an asker and just provide the answer. It's considered rude and won't be tolerated here. 2. Please try and guide the asker through the problem instead of just giving the answer. Thanks.

OpenStudy (anonymous):

Okay, sorry, I'm new here :) By the way, ups I somehow got 8 from 9-2 :D

OpenStudy (anonymous):

lol @addinol